Session: Sep 2000 (session 3)
Wavelength: 18 cm


To give feedback on this experiment, click here

Experiment Feedback

Medicina (Cecchi at Thu Sep 28 07:57:54 METDST 2000) Minor failures
stop recording from 12:20:42 to 12:32:42 W75N because of a tape reel motor failure. strong interference in band

Shanghai (Huang Xinyong at Thu Sep 28 08:10:57 METDST 2000) Success

Torun (K.Borkowski at Thu Sep 28 09:17:23 METDST 2000) Minor failures
As in the previous experiment, the BBC7 and BBC6 malfunctioned. No other known problems.

Noto (C.Trigilio at Fri Sep 29 15:40:47 METDST 2000) Failure
Did not observe because of a failure of the tape recorder (capstan motor) occurred on 22 Sep.

Onsala (John Conway at Fri Sep 29 15:53:32 METDST 2000) Minor failures
Intermittant RFI in RCP , unfortunately occasionally in observing band

Hartebeesthoek (Jonathan Quick at Sun Oct 1 23:44:38 METDST 2000) Minor failures
Missed first 4 minutes of experiment due to setup problems. Failure of 5V sync power supply leads to following problems Scans 271-1627 thru 271-1631: VC04 failed ? Scans 271-1633 thru 271-1732: No data recorded - schedule halted Scans 271-1744 thru 271-1821: VC04 and VC05 intermittently faulty Otherwise no problems.

Effelsberg (A. Kraus at Wed Oct 4 15:52:38 METDST 2000) Minor failures
No data between UT0412-0430 (technical problems).

Jodrell_Bank (Alastair Gunn at Thu Oct 5 10:01:10 METDST 2000) Minor failures
Occasional momentary failures of telescope servos. May be short periods of missing data. Some slew times were not long enough for the Lovell so some scans are missed.